# Generalized Lotka-Volterra model with sparse interactions: non-Gaussian effects and topological multiple-equilibria phase

### v1

@@ III Non-Gaussianity in the single-equilibrium phase
The results for $P^{(+)}(n,m)$ and $P^{(-)}(n,m)$ in the case $\hat{\sigma}=0.20$ are shown in Fig. 5: it is clear that competitive (positive sign) interactions lead to anti-correlations, while mutualistic (negative sign) interactions induce positive correlations.
@@
Figure 5: Connected correlation functions for exclusively competitive $P^{(+)}(n,m)-\eta(n)\eta(m)$ (top) and exclusively mutualistic $P^{(-)}(n,m)-\eta(n)\eta(m)$ (bottom) interactions, for $\hat{\sigma}=0.20$.
III.1 Absence of a glass phase for a small competition strength $\hat{\mu}$
@@
We recall that in the case of a fully-connected network with symmetric interactions [2], a multiple-equilibria phase is found at $T>0$ for any positive $\mu$, provided that $\sigma$ is large enough. Further increasing $\sigma$ leads then to the emergence of an unbounded growth phase, where species grow indefinitely. On the contrary, in the sparse network we find that for finite temperatures for $\hat{\mu}=0.1$, by increasing $\hat{\sigma}$ the system passes directly from the single equilibrium phase to the unbounded growth one, without any signature of an intermediate glassy phase. An investigation of the phase diagram for different (larger) values of $\hat{\mu}$ is presented in Sec. IV.
@@
At this point, it is necessary to clarify what is meant by unbounded growth phase in the context of sparse networks. While in fully-connected networks the identification of the unbounded growth phase is straightforward, since all species behave identically at equilibrium, the same is not true for sparse networks: here species which undergo unbounded growth can coexist with species which do not. Therefore we define the unbounded growth phase of the gLV model on a sparse network as the phase where at least one species shows unbounded growth (details on the numerical identification of the unbounded growth are reported in the appendix). According to our data, for $\hat{\mu}=0.1$ and $T=0.5,0.75,1.0$, when $\hat{\sigma}$ is increased the system passes directly from single equilibrium to the unbounded growth phase. In Fig. 6 the behaviour of the fraction $N_{UG}/N$ as a function of $\hat{\sigma}$ is plotted, where $N_{UG}$ is the number of species with unbounded growth and $N$ is the total number of species.
@@
Figure 6: Fraction of species that grow indefinitely, for three different temperatures, for $\hat{\mu}=0.1$ and $N=128$. Around $\hat{\sigma}\sim 0.35$ the fraction becomes finite (it deviates from the dashed grey line that represents $N_{UG}=0$) and the unbounded growth phase appears.
@@
In order to provide further evidence of the consistency between our results and others from the literature we have also computed, in the single equilibrium phase, the survival probability $\Phi$, i.e., the ratio between the number of surviving species and the total number of species, and the average population abundance $M=\sum_{i=1}^{N}M_{i}$, as a function of the disorder strength $\hat{\sigma}$. Data for $\Phi$ and $M$ are shown respectively in the upper and the lower panel of Fig. 7. In both panels the behaviour of the corresponding observable is represented up to the value of $\hat{\sigma}$ at which unbounded growth is found (see Fig. 6).
@@
Figure 7: We show, for $\hat{\mu}=0.1$ and $N=128$, and three different temperatures, the survival probability $\Phi$ and the mean abundance $M$, varying $\hat{\sigma}$ up to the appearance of the unbounded growth phase.
IV Multiple equilibria at zero disorder: topological glass phase
@@ IV.1 Thermodynamics from Belief Propagation equations
Studying the equilibrium thermodynamics by means of the single-equilibrium (or replica-symmetric) Belief-Propagation equations for the cavity marginals, Eq. (II) (See Appendix A for details on the algorithm), the criterion for detecting the emergence of a multiple-equilibria/glassy phase is provided by the lack of convergence of the above algorithm. This failure to converge means that the equations derived by assuming the existence of a single equilibrium state are not appropriate to describe the equilibrium thermodynamics of the system and more complicated replica-symmetry-breaking ansatzs (multiple-equilibria adapted assumptions on the equilibrium distribution) are needed to study the cavity marginals [21]. On the basis of this criterion we have computed, for the zero disorder case, $\hat{\sigma}=0$, the critical line in the plane $(\hat{\mu},T)$ which separates the single equilibrium phase from a multiple-equilibria one: data are shown in Fig. 8  [^1] It could be possible that a one-step replica symmetry breaking phase arises on top of the replica symmetric one. In this case, even if the replica symmetric phase is still stable and BP equations converge, in the same region of parameters there could exist an additional stable glassy state. Even if we cannot exclude this situation, we see that, on the left of the critical line in Fig. 8 the Langevin dynamics converges to the unique solution found by BP, a strong evidence that no other glassy state is thermodynamically relevant. The fact that the BP equations do not converge anymore beyond the critical line is a strong indication of the fact that the topologically glassy phase is of the full-replica-symmetry breaking type [21].. We highlight that a linear extrapolation of the transition line down to $T=0$ nicely matches the critical value $\hat{\mu}_{c}$ found in [17]. The extrapolation is needed because eqs. (II) are only well defined for $T\neq 0$. We stress that the multiple equilibria phase found here upon increasing $\hat{\mu}$ at zero disorder is quite different from the multiple equilibria phase found upon increasing $\sigma$ at finite $T$ in the fully-connected case  [2]. In the sparse network the presence of many equilibria is determined by the following mechanism: when $\hat{\mu}>\hat{\mu}_{c}(T)$ the interactions between species become too competitive and this causes the extinction of some species. For even higher values of $\hat{\mu}$ the graph breaks into disconnected groups of surviving species. Since all species are equivalent—having the same number of neighbours and identical interactions—the extinction of specific species is solely determined by the randomness associated to the initial values of the marginals and to thermal fluctuations. In particular, different instances of the thermal noise can lead to the extinction of different subsets of species in the sparse network, giving rise to a glass phase which we have termed topological, with no counterpart on dense networks. In Fig. 8 the critical line is plotted in the $(\hat{\mu},T)$ plane: the critical value $\mu_{c}(T)$ decreases by increasing the temperature. This result can be easily understood as thermal fluctuations enhance the extinction of some species, shifting $\hat{\mu}_{c}$ further to the left as temperature increases.
@@
Figure 8: Transition line from single to multiple equilibria in the $(\hat{\mu},T)$ space, with $\hat{\sigma}=0$. The transition line has been computed for $N=256$.
@@ IV.2 Thermodynamics from Langevin Dynamics
Figure 9: In the center, the same $(\hat{\mu},T)$ diagram of Fig. 8 is reported. In this plot we identify six specific values of $(\hat{\mu},T)$ (identified by different markers), for which we show a single dynamics for the single-species abundance populations of a single ecosystem. The multiple equilibria phase appears when extinctions appear. In the plots above, for each species $i$, at each time $t$, $n_{i}(t)$ is the average over the previous $250$ time-steps. Here $N=256$.
@@
The results of Langevin dynamics are in agreement with the results obtained from the Belief Propagation equations. In the phase which, according to BP, is of single-equilibrium, the Langevin dynamics produces trajectories with small fluctuations and no extinctions. On the contrary, as soon as one tries to run the Langevin dynamics in the multiple-equilibria phase, an abrupt spreading of $n_{i}(t)$ trajectories is found, accompanied by the immediate appearing of extinctions. This peculiar behaviour is shown in the panels of Fig. 9: data are taken from the simulations of a random regular graph with $N=256$ species and the value of the parameters $\hat{\mu}$ and $T$ corresponding to each panel is indicated with a symbol (different in shape and color) in the ($\hat{\mu}$,$T$) phase diagram in the center.
@@
Let us comment the simulations of dynamics in Eq. (21) by starting from the $T=0$ case. By looking at the panels at the bottom of Fig. 9 corresponding to $T=0$ and $\hat{\mu}=0.35,0.353$, it is clear that Langevin dynamics converges to the same abundance value for every species. On the contrary, the bottom-right panel of Fig. 9, corresponding to $T=0$ and $\hat{\mu}=0.358$, clearly shows the appearance of extinctions and the spreading of $n_{i}(t)$ trajectories, consistently with the nature of the topological glass phase discussed above. The same transition from a single-equilibrium to a topological multiple-equilibria phase is found by increasing the temperature at fixed $\hat{\mu}$, provided we choose a value $\hat{\mu}<\hat{\mu}_{c}(T=0)$, i.e., smaller than the critical one at zero temperature. This is clearly shown by the two top-right panels, corresponding to ($\hat{\mu}=0.35$, $T=0.6$) and to ($\hat{\mu}=0.353$, $T=0.2$), which show a spreading of $n_{i}(t)$ and the appearance of extinctions, at variance with the behaviour of trajectories at the same values of $\hat{\mu}$ and $T=0$. All panels in Fig. 9 represent trajectories of abundances $n_{i}(t)$ obtained at different values of $\hat{\mu}$ and temperature $T$ by exploiting the same initial conditions and stochastic noise realization. Clearly, different initial conditions and different realization of the noise would have led to the extinction of different species. Finally, by comparing the top-left panel ($\hat{\mu}=0.33$, $T=0.6$) to the top-central one ($\hat{\mu}=0.35$, $T=0.6$) of Fig. 9, it is clear how the same transition from single to multiple equilibria can be found by increasing $\hat{\mu}$ at fixed temperature.
@@
While the results of Langevin dynamics are perfectly compatible with the results obtained from the Belief Propagation equations, we want to stress that from the Langevin approach the transition line between single and multiple equilibria cannot be defined as sharply as from the lack of convergence of BP, which turns out to be a much more powerful tool to study this phase diagram.
@@ IV.3 A re-entrant transition
We have shown that in the case of a sparse interaction network it is possible to find a topological glass phase, characterized by the presence of extinctions. This phase, in contrast with the glass phase typically found in fully-connected models, is due to highly competitive interactions rather than disordered ones. To fully characterize the transition line in the plane $(\hat{\mu},T)$ at $\sigma=0$, we performed a systematic study of the system as the temperature is varied for four fixed values of the (ordered) interaction strength $\hat{\mu}=0.1,0.15,0.2,0.3$. At $\hat{\mu}=0.1$ we are able to heat up the system without ever losing convergence of the algorithm, a signal that no glassy phase is encountered. Instead, at $\hat{\mu}=0.15,0.2,0.3$, we find a critical temperature that divides a single-equilibria phase from a multiple equilibria one increasing $T$ from low temperature, $T_{c}^{\text{down}}(\hat{\mu})$ (black dots in Fig. 10), and another critical temperature that divides a single-equilibria phase from a multiple equilibria one decreasing $T$ from high temperature, $T_{c}^{\text{up}}(\hat{\mu})$ (red dots in Fig. 10). These findings suggest the following scenario: there is a critical line enclosing the topological glass phase which starts at $\hat{\mu}_{c}$ at $T=0$, moves at smaller $\hat{\mu}$ for increasing $T$ and then bends to the right at a critical value of the interactions strength $\mu^{*}$ such that $0.150.1<\mu^{*}<$ 0.15 and goes to some horizontal asymptote at a temperature smaller than $T=15$, for which we checked that the system is in the replica symmetric phase for any $\hat{\mu}$. The re-entrant transition in the $(\hat{\mu},T)$ plane is pictured as a gray dashed line in Fig. 10. A more refined study of the phase diagram will be discussed elsewhere.
From a general perspective, the critical line sketched in Fig. 8 represents an ergodicity-breaking transition remarkably different from the standard replica-symmetry-breaking transition found in fully-connected graphs. In the latter case the transition from single to multiple equilibria is always triggered by the decrease of thermal fluctuations. In contrast, the sparse gLV model exhibits the opposite behavior. The reentrant transition of Fig. 10 is very similar to the inverse freezing transition, present in the Random Blume-Capel model [25, 26]. It is known that, in order to have an inverse-freezing transition, it is essential the presence of active and inactive nodes, that are modeled as bosonic spins that can take values $(+1,-1,0)$ in the Blume-Capel model, while in the sparse gLV the active and inactive nodes correspond to surviving and extincted species.
Figure 10: Transition line, at $\hat{\sigma}=0$, between the single equilibrium and the multiple equilibria phase. The blue line reproduces the transition line from Fig. 8. The black dots are the transition points $T_{c}^{\text{down}}$ found by increasing the temperature from below, while the red dots correspond the transition points $T_{c}^{\text{up}}$ found by decreasing the temperature from above. Based on these points, we suggest that the transition line eventually bends to the right, forming a nose-like shape, which we represent as the grey dashed line in the figure.

### v2

@@ III Strong Non-Gaussianity in the single-equilibrium phase
The results for $P^{(+)}(n,m)$ and $P^{(-)}(n,m)$ in the case $\hat{\sigma}=0.20$ are shown in Fig. 5: it is clear that competitive (positive sign) interactions lead to anti-correlations, while mutualistic (negative sign) interactions induce positive correlations. From the joint probabilities $P^{(+)}(n,m)$ and $P^{(-)}(n,m)$, we can extract the mean values $\langle nm\rangle_{(+)}$ and $\langle nm\rangle_{(-)}$, which correspond to the average product $nm$ for neighbouring species with positive and negative interactions, respectively. Figure 6 shows the corresponding normalized covariances defined as $\left(\langle nm\rangle_{(+)/(-)}-\langle n\rangle\langle m\rangle\right)/\langle n\rangle\langle m\rangle$, plotted as function of $\hat{\sigma}$. Competitive interactions lead to negative correlations while mutualistic interactions to positive correlations. The magnitude of these correlations increases with $\hat{\sigma}$.
@@
Figure 5: Connected correlation functions for exclusively competitive $P^{(+)}(n,m)-\eta(n)\eta(m)$ (top) and exclusively mutualistic $P^{(-)}(n,m)-\eta(n)\eta(m)$ (bottom) interactions. We are inside the strong-nonGaussian region $\hat{\sigma}=0.20$.
Figure 6: Normalized covariances of the average abundances for neighboring species with positive (blue dots) and negative (red dots) interactions. Data are displayed in function of $\hat{\sigma}$. Notice that for low values of $\hat{\sigma}$ no red dots are present because interactions are never mutualistic for low values of $\hat{\sigma}$. Correlations between neighboring species grow (in modulus) increasing the variance of the interactions $\hat{\sigma}$. Error bars are included in the plot.
III.1 Absence of a multiple attractor phase for a small competition strength $\hat{\mu}$
@@
We recall that in the case of a fully-connected network with symmetric interactions [1], a multiple-equilibria phase is found at $T>0$ for any positive $\mu$, provided that $\sigma$ is large enough. Further increasing $\sigma$ leads then to the emergence of an unbounded growth phase, where species grow indefinitely. On the contrary, in the sparse network we find that for finite temperatures and $\hat{\mu}=0.1$, by increasing $\hat{\sigma}$ the system passes directly from the single equilibrium phase to the unbounded growth one, without any signature of an intermediate multiple attractor phase. The convergence of the belief propagation equations on a random graph directly implies the stability of the replica symmetric solution [32]; We have thus checked the convergence of the BP equations (or the entrance in an unbounded growth phase) in the whole range of temperatures for $\hat{\mu}=0.1$ and this is sufficient to show the stability of the RS phase. This is a crucial difference between the BP method and the replica method usually used in fully-connected models: while in the latter, after having done a replica computation with the RS ansatz, one should always check that the RS solution is stable, for BP equations in the RS ansatz instead, the convergence of BP equations automatically implies that the RS phase found is locally stable. We have also checked this result by running the Langevin dynamics many times from different initial conditions, finding either a single fixed point, or an unbounded growth region without signatures of multiple fixed points, as shown in Appendix D. An investigation of the phase diagram for different (larger) values of $\hat{\mu}$ is presented in Sec. IV.
@@
At this point, it is necessary to clarify what is meant by unbounded growth phase in the context of sparse networks. While in fully-connected networks the identification of the unbounded growth phase is straightforward, since all species behave identically at equilibrium, the same is not true for sparse networks: here species which undergo unbounded growth can coexist with species which do not. Therefore we define the unbounded growth phase of the gLV model on a sparse network as the phase where at least one species shows unbounded growth (details on the numerical identification of the unbounded growth are reported in the appendix C). According to our data, for $\hat{\mu}=0.1$ and $T=0.5,0.75,1.0$, when $\hat{\sigma}$ is increased the system passes directly from single equilibrium to the unbounded growth phase. In Fig. 7 the behaviour of the fraction $N_{UG}/N$ as a function of $\hat{\sigma}$ is plotted, where $N_{UG}$ is the number of species with unbounded growth and $N$ is the total number of species.
@@
Figure 7: Fraction of species that grow indefinitely. Data are displayed for three different temperatures, for $\hat{\mu}=0.1$ and $N=128$. Around $\hat{\sigma}\sim 0.35$ the fraction becomes finite (it deviates from the dashed grey line that represents $N_{UG}=0$) and the unbounded growth phase appears.
@@
In order to provide further evidence of the consistency between our results and others from the literature we have also computed, in the single equilibrium phase, the survival probability $\Phi$, i.e., the ratio between the number of surviving species and the total number of species, and the average population abundance $M=\sum_{i=1}^{N}M_{i}/N$, as a function of the disorder strength $\hat{\sigma}$. Data for $\Phi$ and $M$ are shown respectively in the upper and the lower panel of Fig. 8. In both panels the behaviour of the corresponding observable is represented up to the value of $\hat{\sigma}$ at which unbounded growth is found (see Fig. 7).
@@
Figure 8: Survival probability $\Phi$ (top) and the mean abundance $M$ (bottom) in the unique fixed point phase. The observables are displayed for three different temperatures, varying $\hat{\sigma}$, up to the appearance of the unbounded growth phase. The parameters of the model are $\hat{\mu}=0.1$ and $N=128$.
IV Topological multiple equilibria phase at zero disorder
@@ IV.1 Thermodynamics from Belief Propagation equations
Studying the equilibrium thermodynamics by means of the single-equilibrium (or replica-symmetric) Belief-Propagation equations for the cavity marginals, Eq. (II) (See Appendix A for details on the algorithm), the criterion for detecting the emergence of a multiple-equilibria phase is provided by the lack of convergence of the above algorithm. This failure to converge means that the equations derived by assuming the existence of a single equilibrium state are not appropriate to describe the equilibrium thermodynamics of the system and more complicated replica-symmetry-breaking ansatzs (multiple-equilibria adapted assumptions on the equilibrium distribution) are needed to study the cavity marginals [20]. In Appendix G we study the local stability of the unique fixed point found by BP, and we show that indeed when the BP algorithm stops to converge the unique fixed point becomes unstable. On the basis of this criterion we have computed, for the zero disorder case, $\hat{\sigma}=0$, the critical line in the plane $(\hat{\mu},T)$ which separates the single equilibrium phase from a multiple-equilibria one: data are shown in Fig. 9  [^1] It could be possible that a one-step replica symmetry breaking (1RSB) phase arises on top of the replica symmetric one. In this case, even if the replica symmetric phase is still stable and BP equations converge, in the same region of parameters there could exist an additional stable 1RSB state. Even if we cannot exclude this situation, we see that, on the left of the critical line in Fig. 9 the Langevin dynamics converges to the unique solution found by BP, a strong evidence that no other states are thermodynamically relevant. The fact that the BP equations do not converge anymore, and the unique fixed point looses stability beyond the critical line are strong indications of the fact that the topologically multiple equilibria phase is of the full-replica-symmetry breaking type [20], also called Gardner phase in ref. [1]. A precise study of the multiple attractor phase will be performed in a subsequent paper.. We highlight that a linear extrapolation of the transition line down to $T=0$ nicely matches the critical value $\hat{\mu}_{c}$ found in [16]. The extrapolation is needed because eqs. (II) are only well defined for $T\neq 0$. We stress that the multiple equilibria phase found here upon increasing $\hat{\mu}$ at zero disorder is quite different from the multiple equilibria phase found upon increasing $\sigma$ at finite $T$ in the fully-connected case  [1]. In the sparse network the presence of many equilibria is determined by the following mechanism: when $\hat{\mu}>\hat{\mu}_{c}(T)$ the interactions between species become too competitive and this causes the extinction of some species. For even higher values of $\hat{\mu}$ the graph breaks into disconnected groups of surviving species. Since all species are equivalent—having the same number of neighbours and identical interactions—the extinction of specific species is solely determined by the randomness associated to the initial values of the marginals and to thermal fluctuations. In particular, different instances of the thermal noise can lead to the extinction of different subsets of species in the sparse network, giving rise to a multiple attractor phase which we have termed topological, with no counterpart on dense networks. In Fig. 9 the critical line is plotted in the $(\hat{\mu},T)$ plane: the critical value $\mu_{c}(T)$ decreases by increasing the temperature. This result can be easily understood as thermal fluctuations enhance the extinction of some species, shifting $\hat{\mu}_{c}$ further to the left as temperature increases.
@@
Figure 9: Transition line from single to multiple equilibria in the $(\hat{\mu},T)$ space. The transition line has been computed for zero disorder $\hat{\sigma}=0$ and $N=256$.
@@ IV.2 Thermodynamics from Langevin Dynamics
Figure 10: Comparison between the dynamics and the critical line obtained by BP. In the center, the same $(\hat{\mu},T)$ diagram of Fig. 9 is reported. We identify eight specific values of $(\hat{\mu},T)$ (identified by different markers), for which we show a single dynamics for the single-species abundance populations of a single ecosystem. When the multiple equilibria line is crossed, extinctions appear. For each species $i$, at each time $t$, $n_{i}(t)$ is the average over the previous 250250 time-steps. Here $N=256$.
@@
The results of Langevin dynamics are in agreement with the results obtained from the Belief Propagation equations. In the phase which, according to BP, there is a single-equilibrium, the Langevin dynamics produces trajectories with small fluctuations and no extinctions. On the contrary, as soon as one tries to run the Langevin dynamics in the multiple-equilibria phase, an abrupt spreading of $n_{i}(t)$ trajectories is found, accompanied by the immediate appearing of extinctions. This peculiar behaviour is shown in the panels of Fig. 10: data are taken from the simulations of a random regular graph with $N=256$ species and the value of the parameters $\hat{\mu}$ and $T$ corresponding to each panel is indicated with a symbol (different in shape and color) in the ($\hat{\mu}$,$T$) phase diagram in the center.
@@
Let us comment the simulations of dynamics in Eq. (21) by starting from the $T=0$ case. By looking at the panels at the bottom of Fig. 10 corresponding to $T=0$ and $\hat{\mu}=0.35,\,0.353$, it is clear that Langevin dynamics converges to the same abundance value for every species. On the contrary, the bottom-right panels of Fig. 10, corresponding to $T=0$ and $\hat{\mu}=0.354,\,0.358$, clearly show the appearance of extinctions and the spreading of $n_{i}(t)$ trajectories, consistently with the nature of the topological multiple-equilibria phase discussed above. The same transition from single to multiple equilibria can be found by increasing $\hat{\mu}$ at fixed finite temperature, as can be seen by looking at the differences between the top-left panels ($\hat{\mu}=0.341,\,0.347$, $T=0.4$) and the top-right ones ($\hat{\mu}=0.351,\,0.357$, $T=0.4$) of Fig. 10. Please note that in all the right panels, with parameters inside the multiple-equilibria phase, both at zero and at finite temperature, the dynamics firstly seem to converge for small times towards values of the abundances similar to the ones in the unique fixed point, before the appearance of extinctions and the spreading of $n_{i}(t)$ trajectories at large times, clear indications of the multiple-attractor phase. This first approximate convergence at small times is due to a reminiscence of the unique fixed point that looses progressively stability as $\hat{\mu}$ grows, as shown in detail in Appendix G. For this reason, the time spent in the vicinity of the unique fixed point decreases progressively moving to the right of the critical line.
@@
All panels in Fig. 10 represent trajectories of abundances $n_{i}(t)$ obtained at different values of $\hat{\mu}$ and temperature $T$ by exploiting the same initial conditions and stochastic noise realization. Clearly, different initial conditions and different realization of the noise would have led to the extinction of different species, as we show in Appendix H. While the results of Langevin dynamics are perfectly compatible with the results obtained from the Belief Propagation equations, we want to stress that from the Langevin approach the transition line between single and multiple equilibria cannot be defined as sharply as from the lack of convergence of BP, which turns out to be a much more powerful tool to study this phase diagram.
@@ IV.3 A re-entrant transition
[…]
From a general perspective, the replica-symmetry-breaking critical line sketched in Fig. 9 has a shape remarkably different from the standard replica-symmetry-breaking transition line found in fully-connected graphs [1]. In the latter case, the transition from single to multiple equilibria is always triggered by the decrease of thermal fluctuations. In contrast, the sparse gLV model exhibits the opposite behavior. The reentrant transition of Fig. 11 is very similar to the inverse freezing transition, present in the Random Blume-Capel model [8, 14]. It is known that, in order to have an inverse-freezing transition, it is essential the presence of active and inactive nodes, that are modeled as bosonic spins that can take values (+1,1,0)(+1,-1,0) in the Blume-Capel model, while in the sparse gLV the active and inactive nodes correspond to surviving and extincted species.
Figure 11: Full transition line, at $\hat{\sigma}=0$, between the single equilibrium and the multiple equilibria phase. The blue line reproduces the transition line from Fig. 9. The black dots are the transition points $T_{c}^{\text{down}}$ found by increasing the temperature from below, while the red dots correspond the transition points $T_{c}^{\text{up}}$ found by decreasing the temperature from above. Based on these points, we suggest that the transition line eventually bends to the right, forming a nose-like shape, which we represent as the grey dashed line in the figure.